If a Hot Dog has 150 calories per link and the fat content is 14 grams, how many calories are from fat and what percentage is that? Remember to keep straight what unit you are working with. You may need to change one unit to another (hint, hint). Show your work.
Answer #1

Answer: Solution:

Given;

Fat content in hot dog = 14 grams.

So, 1 gram of fat = 9 calories

Therefore: 14 grams of fat will have = 14 ×9 = 126 calories.

So, the fat content 14 grams will have 126 calories.

Percentage= 126/150 = 0.84 × 100 = 84%.